If a force of 25N is applied on a 10kg mass, it accelerates at 1 m/s^2. Find the co-efficient of force between the block and the surface.
Answer #1

Solution)

We know,

Force, F= us* mg

25 N= us * 10kg*9.8 m/s^2

So, coefficient of force, us=25/(10*9.8) = 0.255

==============